# JaxRobotarium: Training and Deploying Multi-Robot Policies in 10 Minutes

### v2

@@ 8 Conclusion
We contribute JaxRobotarium, an accelerated benchmark platform for MRRL with free, publicly available, and standardized sim-to-real evaluation through the Robotarium hardware test bed. Our benchmark experiments demonstrate up to a 150x speed up in simulation, and a 20x speed up in training multi-robot policies compared to baseline. We demonstrate the utility of our platform by evaluating 4 MARL algorithms on 8 multi-robot coordination scenarios, with a large-scale sim-to-real evaluation consisting of over 200 real world deployments. We hope JaxRobotarium will be a valuable resource for accelerating the development, benchmarking, and sim-to-real evaluation of MRRL algorithms and architectures.
@@ 9 Limitations
We design our platform for sim-to-real evaluation through the Robotarium testbed, so we only simulate the Robotarium GRITSBots and we do not simulate observations from more complex perception modules, such as images or LiDAR scans.
The decision to utilize the Robotarium introduces additional limitations. Since the robots are homogenous, heterogeneity has to be artificially injected, such as limiting a robot’s velocity or waypoint step size, or through introducing scenario-specific heterogeneity that is recognized by the scenario’s step function, such as obeying the foraging level of robots when deciding if a resource is successfully foraged. Furthermore, the Robotarium hardware test bed does not support Jax libraries, so the quadratic problem solver used for barrier certificates in deployment and the solver used in Jax simulation (chosen differently for compatibility with jit and vmap) are not aligned. We observe that underlying differences in the solvers can result in different proposed solutions to the same quadratic program.

### v3

@@ 8 Conclusion
We contribute JaxRobotarium, an accelerated benchmark platform for MRRL with free, publicly available, and standardized sim2real evaluation through the Robotarium hardware test bed. Our benchmark experiments demonstrate up to a 150x speed up in simulation, and a 20x speed up in training multi-robot policies compared to baseline. We demonstrate the utility of our platform by evaluating 4 MARL algorithms on 8 multi-robot coordination scenarios, with a large-scale sim2real evaluation consisting of over 200 real world deployments. We hope JaxRobotarium will be a valuable resource for accelerating the development, benchmarking, and sim2real evaluation of MRRL algorithms and architectures.
@@ 9 Limitations
We design our platform for sim2real evaluation through the Robotarium testbed, so we only simulate the Robotarium GRITSBots and we do not simulate observations from more complex perception modules, such as images or LiDAR scans.
The decision to utilize the Robotarium introduces additional limitations. Since the robots are homogeneous, heterogeneity has to be artificially injected, such as limiting a robot’s velocity or waypoint step size, or through introducing scenario-specific heterogeneity that is recognized by the scenario’s step function, such as obeying the foraging level of robots when deciding if a resource is successfully foraged. Furthermore, the Robotarium hardware test bed does not support Jax libraries, so the quadratic problem solver used for barrier certificates in deployment and the solver used in Jax simulation (chosen differently for compatibility with jit and vmap) are not aligned. We observe that underlying differences in the solvers can result in different proposed solutions to the same quadratic program.
We note that our platform cannot be considered a replacement for training in high-fidelity robotic simulation environments such as IsaacGym [16] and IsaacLab [17]. Rather, JaxRobotarium is designed to be a turn-key platform for simulation-training-deployment tailored for MRRL research, offering native integrations (e.g., JaxMARL), standardized benchmarks (e.g., RWARE adaptations), and a ready-to-use zero-cost sim2real pipeline.